OpenStudy (anonymous):

Solve for x: 2|2x-2|+4=20 Help please!!!

OpenStudy (anonymous):

x=5 i do believe. 2|2(5)-2|+4=20 10-2=8 so 2(8)+4=20 16+4=20

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

Yes Thanks so much for your help.